L^{A}TEX template for a project in numerical methods

D.V. Fedorov Århus University

Resumé This is a L^{A}TEX file which can be used as a template for a project in numerical

methods.

1 Section title

Here is a random equation,

H|ψ⟩ = λ|ψ⟩ ,  (1)

which can be later referred to as equation number (1). Figures can be included e.g. using the graphicx package, see figure (1). Figur 1: Numerical solution of a Schrödinger equation with oscillator potential